# A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status

### v3

@@ 3.1. Model Inputs
Infection-hospitalization rate $H_{r}$. We translate infections to projected hospitalizations using an estimated infection-hospitalization rate $H_{r}$ for unvaccinated, infection-naïve males 18–25 years old. $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization. We emphasize that $H_{r}$ is a sex-specific rate. The FDA provides data in Table 2 of [10] indicating that for the 18–24 age-range $H_{r}$ for males is significantly less that that for females. [^7] The data of Table 1 of [54] is consistent with that of [10, Table 2].
Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection prior to the evaluation period.
@@
Hospitalization risk reduction $HRR_{v}$, $HRR_{pi}$, and $HRR_{h}$. These are the fractions by which the hospitalization rate $H_{r}$ is reduced for populations that are protected respectively by vaccination, prior infection, or both (hybrid protection).
VAM/P hospitalizations $H_{\rm VAM/P}$. Our final model input $H_{\rm VAM/P}$ is the projected number of VAM/P hospitalizations expected to occur in attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males. To vaccinate fully 1 million individuals with an mRNA COVID-vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse-events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations. In Section 2.3, based on data available before 1/1/22, we obtained an estimated VAM/P hospitalization rate of 250 per million 2nd doses of mRNA-1273 given to 18–25-year-old males. In Section 6.7 of Supplement 1, we discuss VAM/P-hospitalization risk among 18–25-year-old males for dose 1 of mRNA-1273, justifying a choice of 18 VAM/P hospitalizations per million 1st doses, which is the 1st-dose rate of the FDA’s BLA-Memo Scenario—compare the first two rows of Table 4 above. Thus, for all our scenarios, we set $H_{\rm VAM/P}=268$. [^8] In Supplement 1 (Appendix 2, Section 2.4), we show that our VAM/P hospitalization-rate estimate of 268 per million full mRNA-1273 vaccinations among 18–25-year-old males is consistent with VAM/P data of three major studies [5, 7, 8] appearing in 2022 but available only after 1/22/22, showing the data suggests a VAM/P hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
@@ 3.2. Combining Model Inputs to Produce Outputs
Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1/1/22 – 5/31/22. Let $U$ denote the expected number of Omicron hospitalizations occurring in this test population over the 5-month evaluation period assuming its members remain unvaccinated throughout the period. Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period. The difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section (thus, our model is deterministic).
@@
Next suppose a fraction $F_{pi}$ of the unvaccinated test population has had a COVID-19 infection prior to the start of the evaluation period. We must account for the protection against infection and hospitalization afforded by prior infection. We accomplish this in two stages, first accounting for protection against infection, then for protection from progression to severe disease for infected individuals. The first stage depends on the model input $E_{pi}$, while the second stage uses $HRR_{pi}$.
So, with fraction $F_{pi}$ of the test population previously infected, $(1-F_{pi})\cdot(1{\rm m})$ are infection-naïve and will respond like the baseline population described above. That is, we expect $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}$ infections leading to $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations from this group. Among the remaining $F_{pi}\cdot(1{\rm m})$ members of the test population with prior infection, we expect $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})$ infections resulting in $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})\cdot H_{r}\cdot(1-HRR_{pi})$ hospitalizations. Thus the total number of hospitalizations expected in the unvaccinated test population is
@@
Deriving $V$. Next we assume that every individual in the test population is fully vaccinated. Since all members of this test population are vaccinated, those with prior infection have hybrid protection. Once again we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals. Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
Then, in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection. So the total number of hospitalizations in the population is
@@
We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent in our model to assuming that all persons in the test population are infection-naïve, i.e. that $F_{pi}=0$. With $F_{pi}=0$ in Equation (3.2.3), its right-hand side is
@@
[…]
@@ 3.3. Estimating Values for Model Inputs
Like the FDA, we account for uncertainties in model inputs by adopting a scenario-based assessment. The FDA’s model-input assumptions for its six scenarios of [11] are summarized in Table 1 of Section 2. Observe that the first three scenarios differ only in their COVID-incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario. Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’s Scenarios 1–3 because our COVID incidence assumptions are derived from the FDA’s corresponding assumptions. Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A-C; also, for Scenario D our values for other inputs represent our best estimates. Thus, Scenario D is our most likely scenario. In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed in Section 2.2, we believe to be the minimum reasonable assumption of prior-infection’s protectiveness relative to vaccination’s protectiveness.
Table 5 below summarizes model inputs with footnotes providing some comments on the estimation of their values—see Supplement 1 for details.
@@
$E_{pi}$: Effectiveness of prior-infection protection against COVID-19 infection $0\leq E_{pi}\leq 1$ 0.47
@@
*Remarks. (i) Estimates are based on data available before 1/22/22. (ii) $E_{pi}$, $E_{v}$, and $E_{h}$ are relative to infection-naïve, unvaccinated persons. (iii) $HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$ are relative to unvaccinated persons experiencing primary infections. (iv) Our Scenario-A value $I_{r}=1$ reflects the assumption that all infection-naïve unvaccinated males 18–25 become infected over the evaluation period, consistent with FDA Scenario-1 assumptions (see Section 2.2). (v) We estimate $H_{r}$ using an IHR model [55] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [56]. We assume a 40% incidental-hospitalization rate and make sex-based adjustments based on [10, Table 2]. (See [Supplement 1, Section 2]) (vi) $F_{pi}=0.62$ is the average of (a) the CDC’s estimate that 54.9% of those 18–49 years old had been infected by COVID-19 before 10/1/21 [25] and (b) an estimate derived in Supplement 1 that 69% of unvaccinated male 18–25 year-olds had been infected by 1/1/2022. (vii) For reasons discussed in Supplement 1, for all scenarios, we use $E_{v}=0.3$ and $HRR_{v}=0.67$; thus, we are assuming a VEH of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the FDA’s VEH (viii) Estimates for $E_{pi},E_{h},HRR_{pi}$, and $HRR_{h}$, are based on data from [26] and [27]. $HRR_{pi}=0.79$ and $HRR_{h}=0.88$ combine with $E_{pi}=0.47$ and $E_{h}=0.57$ to yield overall estimates for protection against hospitalization—88.9% for prior infection and 94% for hybrid protection—consistent with those of meta-analyses [21] and [22].
@@ 4. Reanalysis: Results
* For discussions of “Assumptions,” see Section 3.3 as well as Supplement 1.
We illustrate our model’s application by carrying out Scenario-A computations, Using Scenario A assumptions from Table 6 and our definitions (3.2.1) and (3.2.2) of $U$ and $V$, we obtain
$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.47\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 629,{\rm and}\\$ $V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
Thus, for Scenario A, the hospitalizations-prevented benefit of vaccination is projected to be UV= 629168=461U-V=\ 629-168=461 per million full mRNA-1273 vaccinations, roughly one-tenth of the FDA’s Scenario-1 estimate 4,766 [11]. Although Scenario A has a benefit-risk ratio above 1, its assumption that $I_{r}=1$ makes this scenario highly unlikely. [^9] Based on seroprevalence and modeling information, we estimate $I_{r}\leq 0.5$ [Supplement 2, Section 2.3]. For the more likely scenarios B, D, and E, our modeling suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from VAM/P alone compared to COVID hospitalizations prevented. Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67. Though Scenario D is our most likely scenario, Scenario E is equally important, being a “best-case scenario” for vaccination incorporating the minimum reasonable assumption of prior-infection protection vs. vaccine protection.
In Section 2 of Supplement 2, we carry out sensitivity analyses exploring how modifying model inputs impacts the benefit-risk picture. For example, if we fix our model-input values, except for $I_{r}$, at their Scenario-D levels, then benefit-risk equipoise is reached when $I_{r}\approx 0.685$. We also illustrate how slider-plots can provide insights concerning equipoise questions involving variation of more than one model input.

### v4

@@ 2.1. Model Inputs
Infection-hospitalization rate $H_{r}$. We translate infections to projected hospitalizations using an estimated infection-hospitalization rate $H_{r}$ for unvaccinated, infection-naïve males 18–25 years old. $H_{r}$ is the fraction of a population of unvaccinated individuals experiencing a primary COVID infection that will progress to severe disease requiring hospitalization.
Prior-infection fraction $F_{pi}$. This is the fraction of the test population having had a COVID-19 infection before the start of the evaluation period.
@@
Hospitalization-risk reduction $HRR_{v}$, $HRR_{pi}$, and $HRR_{h}$. These are the fractions by which the hospitalization rate $H_{r}$ is reduced for populations that are protected, respectively, by vaccination, prior infection, or both (hybrid protection).
VAM/P hospitalizations $H_{\rm VAM/P}$. Our final model input $H_{\rm VAM/P}$ is the projected number of VAM/P hospitalizations expected to occur in attaining a test population of 1 million fully mRNA-1273 vaccinated 18–25-year-old males. To fully vaccinate 1 million individuals with an mRNA COVID vaccine, more than 2 million doses must be administered (assuming, e.g., that those who experience myo/pericarditis related to dose 1 do not take dose 2), and the adverse events associated with each of the over 2 million doses reflect the cost to the population of obtaining the benefits of 1 million full vaccinations.
@@ 2.2. Combining Model Inputs to Produce Outputs
Recall that we are considering a test population of 1 million males 18–25 over the evaluation period 1 January 2022–31 May 2022. Let $U$ denote the expected number of Omicron hospitalizations occurring in this test population over the 5-month evaluation period assuming its members remain unvaccinated throughout the period. Let $V$ denote the corresponding number of expected hospitalizations assuming the population is vaccinated throughout the period. The difference UVU-V is the hospitalizations-prevented benefit of vaccination for the test population, and the outputs $U$ and $V$ can be written in terms of the model inputs defined in the preceding section (thus, like the FDA’s, our model is deterministic).
@@
Next, suppose a fraction $F_{pi}$ of the unvaccinated test population has had a COVID-19 infection prior to the start of the evaluation period. We must account for the protection against infection and hospitalization afforded by prior infection. We accomplish this in two stages, first accounting for protection against infection, then for protection from progression to severe disease for infected individuals. The first stage depends on the model input $E_{pi}$, while the second stage uses $HRR_{pi}$.
So, with fraction $F_{pi}$ of the test population previously infected, $(1-F_{pi})\cdot(1{\rm m})$ are infection-naïve and will respond like the baseline population described above. That is, we expect $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}$ infections leading to $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot H_{r}$ hospitalizations from this group. Among the remaining $F_{pi}\cdot(1{\rm m})$ members of the test population with prior infection, we expect $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})$ infections resulting in $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{pi})\cdot H_{r}\cdot(1-HRR_{pi})$ hospitalizations. Thus, the total number of hospitalizations expected in the unvaccinated test population is
@@
Deriving $V$. Next, we assume that every individual in the test population is fully vaccinated. Since all members are vaccinated, those with prior infection have hybrid protection. Once again, we reason in two stages, first accounting for the protection against infection, then for the protection against hospitalization for infected individuals. Once again, we assume a fraction $F_{pi}$ of the test population has had a prior COVID-19 infection.
Then, in the fully vaccinated test population, the expected number of hospitalizations is $(1-F_{pi})\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{v})\cdot H_{r}\cdot(1-HRR_{v})$ for those without prior infection and $F_{pi}\cdot(1{\rm m})\cdot I_{r}\cdot(1-E_{h})\cdot H_{r}\cdot(1-HRR_{h})$ for those with hybrid protection. So, the total number of hospitalizations in the population is
@@
We remark that disregarding the protection afforded by prior infection, as the FDA did, is equivalent in our model to assuming that all persons in the test population are infection-naïve, i.e., that $F_{pi}=0$. With $F_{pi}=0$ in Equation (2.2.3), its right-hand side is
@@
which is essentially the FDA’s model $H\cdot VEH$ for hospitalizations prevented. The quantity $1-(1-E_{v})\cdot(1-HRR_{v})$ of (4) corresponds to the $VEH$ formula on page 10 of [54], which is the reference the FDA used in [13] for its $VEH$ estimate of 0.72. Thus, our model for hospitalizations-prevented UVU-V represents a natural extension of the FDA’s model.
@@ 2.3. Estimating Values for Model Inputs
Like the FDA, we account for uncertainties in model inputs by adopting a scenario-based assessment. The FDA’s model-input assumptions for its six scenarios of [13] are summarized in Table 1 above. Observe that the first three scenarios differ only in their COVID incidence assumptions, with Scenario 1 representing the FDA’s most likely scenario. Our first three scenarios, Scenarios A–C, are reanalyses, respectively, of the FDA’s Scenarios 1–3 because our COVID incidence assumptions are derived from the FDA’s. Based on information available at the time of the FDA’s mRNA-1273 analysis, we consider the COVID incidence assumption of our Scenario D more likely to be accurate than those of Scenarios A–C; also, for Scenario D, our values for other inputs represent our best estimates. Thus, Scenario D is our most likely scenario. In Scenario E, our final scenario, we assume that prior infection and mRNA-1273 vaccination provide equivalent protection against Omicron, which, for reasons discussed Section 1.3, we believe to be the minimum reasonable assumption of prior infection’s protectiveness relative to vaccination’s protectiveness. Model inputs and outputs for our scenarios are described in the next section, while the methods we use to estimate input values are described below and summarized in Table 4. Our estimates are based on data available to the FDA before 22 January 2022.
Estimating $I_{r}$. As we noted in Section 1.3, the FDA’s COVID-case-rate assumption for its most likely scenario, Scenario 1, corresponds to the assumption that essentially all unprotected (i.e., unvaccinated) 18–25-year-old males become infected over the 5-month evaluation period. Thus, in Scenario A, our reanalysis of Scenario 1, we too assume all unprotected persons in the test population become infected. However, in our modeling, the unprotected are the infection-naïve unvaccinated. Assuming all such persons become infected over 5 months is equivalent to setting $I_{r}=1$, our Scenario-A value for $I_{r}$. As discussed in Section S1 of Supplement S1, for our Scenarios B and C, values of $I_{r}$ are determined by the FDA’s COVID-incidence assumptions for its Scenarios 2 and 3, respectively, while for our Scenarios D and E, $I_{r}$ is derived from the assumption that over the evaluation period the increase in COVID-infection level among 18–25-year-old males is twice that of the general population during the second COVID wave in the U.S. (1 October 2020–28 February 2021).
Estimating $H_{r}$. We estimate $H_{r}$ using an IHR model [55] developed via a seroprevalence-based meta-analysis anchored to the ancestral strain of SARS-CoV-2 and available as a preprint in October 2021 [39]. Applying the model, we obtain a “with COVID” IHR of 0.37% for the infection-naïve unvaccinated of ages 18–25 and reduce this rate assuming 40% of hospitalizations attributed to COVID are incidental (see Section 1.3), arriving at an IHR of $0.22\%\approx 0.6\cdot 0.37\%$ (for infection-naïve unvaccinated 18–25-year-olds). The 0.22% rate for the general population of 18–25-year-olds in the U.S. (males and females) is consistent with U.S. military COVID-hospitalization data [56]. See Supplement S1 for further details concerning our derivation of the IHR value $0.22\%$.
In its benefit–risk assessments [12, 13] for, respectively, Pfizer’s and Moderna’s mRNA vaccines, the FDA relied on sex-specific CDC hospitalizations data. The case and hospitalization rates per 100,000 persons in Table 2 of [12] suggest that for the 18–24 age range, females are 2.53 times as likely as males to be hospitalized with COVID-19 and 1.21 times as likely to have a case of COVID-19.888These relationships are consistent with those of [57, Table 1] based on data from 1 March 2020 to 1 March 2022. In Section S2 of Supplement S1, we use hospitalization and case data from Table 2 of [12] to estimate the IHR for unvaccinated, infection-naïve males 18–25 to be $0.14\%$; thus, we set $H_{r}=0.0014$. This value agrees with a natural IHR estimate derived from the FDA’s BLA-Memo scenario (discussed in Section 1.5). For the BLA-Memo Scenario, vaccination prevents 76,362 cases and 1755 hospitalizations by providing 30% effectiveness against cases and 72% effectiveness against hospitalizations. Thus, the scenario assumes that $76,362/0.3=254,540$ cases among the unvaccinated would lead to $1755/0.72\approx 2438$ hospitalizations, yielding a case hospitalization rate of $\approx$0.958%. Applying the CDC case-to-infection multiplier of 4 and assuming a 40% incidental COVID-hospitalization rate, we arrive at the estimate $H_{r}=0.6\cdot(2438)/(4\cdot$ 254,540) $\approx$ 0.0014.
Estimating $F_{pi}$. For our Scenarios A–C, we use the value $F_{pi}=0.62$, which is the average of (a) the CDC’s estimate, mentioned earlier, that 54.9% of those 18–49 years old had been infected by COVID-19 before 1 October 2021 [31] with (b) an estimate derived in Supplement S1 that 69% of unvaccinated male 18–25-year-olds had been infected by 1 January 2022. For Scenarios D and E, we use $F_{pi}=0.69$. These values of $F_{pi}$ may be underestimates because they rely on the CDC’s infected rate of 54.9% by 1 October 2021 for the general population of 18–49 year-olds, and we would expect a somewhat higher infected rate among the unvaccinated.
Estimating $E_{v}$, $HRR_{v}$. Recall that the FDA used $VE=0.30$ and $VEH=0.72$ for all its Omicron-based scenarios [13, Table 1]. For reasons discussed in Supplement S1, for all our scenarios, we use $E_{v}=0.30$ and $HRR_{v}=0.67$. Hence, we are assuming a $VEH$ of $1-(1-0.30)(1-0.67)\approx 0.77$, which exceeds the $VEH$ used by FDA.
Estimating $E_{pi}$ and $E_{h}$. Our values $E_{pi}=0.45$ and $E_{h}=0.57$ are derived (see Supplement S1) from a December 2021 study [32], based on data from South Africa. [^9] As discussed in Supplement S1, our effectiveness estimates involving prior infection are consistent with those of meta-analyses [27] and [28].
Estimating $HRR_{pi}$ and $HRR_{h}$. In Supplement S1, we use data from [33] (based on Omicron cases in England for the period 1 December 2021–14 December 2021) to estimate $HRR_{pi}=0.79$ and $HRR_{h}=0.88$. [^10] The estimates $HRR_{pi}=0.79$ and $HRR_{h}=0.88$ combined with our estimates $E_{pi}=0.45$ and $E_{h}=0.57$ yield overall estimates for protection against hospitalization—88.5% for prior infection and 94% for hybrid protection—consistent with those of meta-analyses [27] and [28]. We note that hybrid protection’s exceeding vaccine-only protection is consistent with studies such as [58, 59]. .
Estimating $H_{\rm VAM/P}$. In Section 1.4, we combined four assessments of VAM/P risk based on data available before 1 January 2022 to obtain an estimate of 250 VAM/P hospitalizations per million 2nd doses of mRNA-1273 given to 18–25-year-old males. In Section S6.7 of Supplement S1, we discuss VAM/P-hospitalization risk among 18–25-year-old males for dose 1 of mRNA-1273, justifying a choice of 18 VAM/P hospitalizations per million 1st doses, which is the 1st-dose risk the FDA uses in its BLA-Memo Scenario—compare the first two rows of Table 3 above. Thus, for all our scenarios, we set $H_{\rm VAM/P}=268$. [^11] In Supplement S1 (Appendix S2, Section S2.4), we show that our VAM/P hospitalization-rate estimate of 268 per million full mRNA-1273 vaccinations among 18–25-year-old males is consistent with VAM/P data of three major studies [5, 7, 8] appearing in 2022 but available only after 22 January 2022, showing the data suggests a VAM/P-hospitalization rate between 240 and 277 per million full, homologous mRNA-1273 vaccinations among males 18–25.
@@
$E_{pi}$: Effectiveness of prior-infection protection against COVID-19 infection $0\leq E_{pi}\leq 1$ 0.45
@@
* Remarks. All measures of effectiveness of protection against infection ($E_{pi}$, $E_{v}$, and $E_{h}$) are relative to infection-naïve, unvaccinated persons. All hospitalization-risk reductions ($HRR_{pi}$, $HRR_{v}$, and $HRR_{h}$) are relative to unvaccinated persons experiencing primary infections.
@@ 3. Reanalysis: Results
* For discussions of “Assumptions,” see Section 2.3 as well as Supplement S1.
We illustrate our model’s application by carrying out Scenario-A computations. Using Scenario A assumptions from Table 5 and our definitions (1) and (2) of $U$ and $V$, we obtain
$U=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot 0.0014+0.62\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.45\right)\cdot 0.0014\cdot\left(1-0.79\right)\approx 632,{\rm and}\\$ $V=\left(1-0.62\right)\cdot\left(1{\rm m}\right)\cdot 1\cdot\left(1-0.30\right)\cdot 0.0014\cdot\left(1-0.67\right)+0.62\cdot 1\cdot\left(1-0.57\right)\cdot 0.0014\cdot\left(1-0.88\right)\approx 168.$
Thus, for Scenario A, the hospitalizations-prevented benefit of vaccination is projected to be UV= 632168=464U-V=\ 632-168=464 per million full mRNA-1273 vaccinations, roughly one-tenth of the FDA’s Scenario-1 estimate 4766 [13]. Although Scenario A has a benefit–risk ratio above 1, its assumption that $I_{r}=1$ makes this scenario highly unlikely. [^12] Based on seroprevalence and modeling information, we estimate $I_{r}\leq 0.5$ [Supplement S2, Section S2.3]. For the more likely scenarios B, D, and E, our modeling suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 62% more hospitalizations for VAM/P alone compared to COVID hospitalizations prevented. To be conservative, we adjust these estimates downward.
As discussed in Section 1.4, for our main analysis (whose results are presented in Table 5 above), we have included 1st dose VAM/P risk without accounting for 1st dose benefits (which is the choice the FDA made in its BNT162b2 assessment [12] and its BLA-Memo assessment of mRNA-1273 [51]). This may be viewed as a conservative choice, being consistent with the principle that vaccines should have “highly favorable benefit risk profiles” [1]. Nevertheless, as we indicated in Section 1.4, including 1st dose VAM/P risk while not including 1st dose benefits potentially biases an assessment against vaccination. Thus, we adjust our results to remove this bias by comparing the hospitalizations-prevented benefit of two doses to the VAM/P risk attributable to dose two only. Although this adjustment potentially biases our outcomes analysis in favor of vaccination, we, conservatively, characterize the outcomes of our more likely scenarios in the following way: comparing our estimated 2nd-dose VAM/P risk at 250 hospitalizations per million to the hospitalizations-prevented benefit of two doses—165 per million in Scenarios B, 181 per million in Scenario D, and 232 per million in Scenario E—we conclude that mRNA-1273 vaccination of 18–25-year-old males generated between $8\%$ and $52\%$ more hospitalizations for VAM/P compared to COVID hospitalizations prevented, with an increase of $38\%$ for Scenario D. In particular, relative to hospitalizations, risks exceed benefits in both Scenario D, our most likely scenario, and in Scenario E, a “best-case scenario” for vaccination incorporating the minimum reasonable assumption of prior-infection protection vs. vaccine protection.
An ideal benefit–risk assessment of a two-dose vaccination series would account for the benefits and risks of both doses. We present such an assessment in Section S2 of Supplement S2 and provide some commentary on it in Section 3.3 below, where we discuss sensitivity analyses exploring how modifying model inputs impacts the benefit–risk picture.